OpenStudy (anonymous):

will fan and give medal state the property that justifies the statement: If 3x + 4 = 10, then 3x = 6 (1 point)

OpenStudy (doc.brown):

This question doesn't really demonstrate any particular Algebraic Property that I have written down, 1. Closure: a + b and ab are real numbers 2. Commutative: a + b = b + a, ab = ba 3. Associative: (a+b) + c = a + (b+c), (ab)c = a(bc) 4. Distributive: (a+b)c = ac+bc 5. Identity: a+0 = 0+a = a 6. Inverse: a + (-a) = 0, a(1/a) = 1 7. Cancelation: If a+x=a+y, then x=y 8. Zero-factor: a0 = 0a = 0 9. Negation: -(-a) = a, (-a)b= a(-b) = -(ab), (-a)(-b) = ab

OpenStudy (doc.brown):

It's really just an example of the golden rule; Any mathematical operation can be performed on one side of an equation so long as the identical operation is performed on the other side of the equation.

OpenStudy (anonymous):

you want the answe choices?

OpenStudy (doc.brown):

Sure

OpenStudy (anonymous):

Subtraction Property of Equality Addition Property of Equality Division Property of Equality Multiplication Property of Equality

OpenStudy (doc.brown):

Ah, I see. They're asking, how did we get rid of 4 from each side?

OpenStudy (anonymous):

i have no clue.

OpenStudy (anonymous):

subtraction... omg im stupid

OpenStudy (doc.brown):

How do you turn 3x+4 into 3x?

OpenStudy (anonymous):

you subtract four from 10 right??

OpenStudy (doc.brown):

Right!

OpenStudy (anonymous):

and then you divide 6 from 3 and get 2 but the answer is the subtracting one right?

OpenStudy (doc.brown):

In the step that gets you from 3x+4 = 10 to 3x=6 yes, it's subtraction

OpenStudy (anonymous):

will you please help me with another one??

OpenStudy (doc.brown):

Sure, just please post it as a new question.

OpenStudy (anonymous):

kk.